OpenStudy (anonymous):

the sum of two numbers is 110. Their difference is 28. what is the smallest of the two numbers? a.56 b.41 c.69 d.55

OpenStudy (anonymous):

41?

OpenStudy (anonymous):

The way to work out this question is to divide 110 by 2, and divide 28 by 2. You get 55 and 14. to get the smaller of the two numbers, take 14 away from 55 which will give you 41 To get the bigger of the 2 add 14 and you get 69 :)

OpenStudy (campbell_st):

set up the 2 equations a + b = 110 (1) a - b = 28 (2) then (1) + (2) will give 2a = 138 a = 69 substitute into (1) b = 41 the smaller number is 41